# ROADWork: A Dataset and Benchmark for Learning to Recognize, Observe, Analyze and Drive Through Work Zones

### v2

@@ B.1 Image Acquisition
Visual data were acquired from cameras mounted inside a vehicle while driving through 18 US cities, resulting in 9650 images from three sources: (a) images that we captured in Pittsburgh (b) images that were semi-automatically extracted from the Michelin Mobility Intelligence (MMI) Open Dataset (formerly RoadBotics) [41] (c) Images that were discovered in Mapillary [74], BDD100K [107] and other other data sources by our models trained on data from the first two sources.
Main Data Sources. To collect the first data subset of the main dataset, we drove on urban, suburban, and rural roads in Pittsburgh and captured 2,338 ($32\%$) images with an iPhone 14 Pro Max paired with a Bluetooth remote trigger. Next, images from other U.S. cities were sourced from videos in the MMI Open Dataset. A combination of Detic [115] and a cone detector trained on NuScenes [5] were used to mine frames presumed to contain roadwork zones with detector confidence at $25\%$ – ensuring high recall with the expense of low precision. This process yielded approximately 100000 candidate images. We then manually selected 5078 (68%) images containing unique road objects or roadwork zones, prioritizing individual scene diversity. The distribution of images across U.S. cities is shown in Figure 12.
Discovered Data Sources. In Section 3 of the main manuscript, we described our model and the work zone classification rule that we use to discover images. We discovered work zone images from common driving datasets (a) 558 images from Mappilary [74] and (b) 411 images from BDD100K [107]. Additionally, we exploit other data sources to curate 1265 images into various subsets containing work zones (See Figure 13 for examples). These subsets were further manually filtered to remove redundant images. We describe the subsets below,
@@
Figure 13: ROADWork: Additional In-The-Wild Discovered and Annotated Work Zone Images. Following the discovery process described in Section 3 of the main manuscript, we discovered and annotated an additional 1265 work zones images from a variety of sources apart from 969 images discovered in BDD100K [107] and Mapillary [74]. (a) The top row depicts Vehicle-Pittsburgh subset images we discovered from driving in Pittsburgh (around 157 images). The subset consists of work zone images taken in bad weather and night. (b) The middle row depicts Vehicle-Rural subset images we driving on rural areas and highways in the US (308 images). The subset consists of work zone images taken in both day and night. (c) The bottom row depicts images discovered from a Bus that was driven on a fixed route in Pittsburgh. We discovered 272 images captured from the front camera, while 528 images were captured from other cameras mounted on the bus. Images were captured in all conditions, including bad weather and night.
@@ B.2 Annotations
The object annotation workflow combined automatic and manual labeling, followed by manual verification. To reduce annotation effort, we used Detic [115] with a custom vocabulary of “cone, drum, vehicle, traffic sign” to bootstrap annotations on our captured images. However, category predictions from Detic [115] were discarded as due to frequent classification errors. Polygons were simplified using the Vishwalingam-Wyatt algorithm [95] to facilitate editing. All object categories were manually assigned, and any additional objects in these images, as well as objects in all other images, were manually segmented and categorized. Finally, all annotations were manually verified by one person.
@@ B.3 Metric 3D Reconstruction and Pathway Generation from Smartphone Videos
Leveraging Smartphone-As-Dashcam Videos. Our work utilizes the MMI Open Data Set [41], which contains extensive video footage captured from a Samsung Galaxy S9 smartphone, for which camera intrinsics are known. From this dataset, we extract 30-second video snippets corresponding to our annotated work zones. These snippets are then downsampled to 5 FPS to yield the final set of smartphone images for our 3D reconstruction pipeline.
@@
Leverging 3D Reconstruction As Anchor. Our primary goal is to produce an accurate, metric-scale, and spatially-aligned 3D reconstruction from the collection of smartphone videos, which have weakly-aligned GPS metadata. With recent advances in Visual Place Recognition [1], it’s likely that the weakly-aligned GPS metadata might also be superfluous in the future.
@@
The core of our approach is to anchor our reconstruction to a set of images with high-quality pose information [97, 98]. To achieve this, we use the initial, coarse GPS from each video to query and retrieve nearby Google Street View panoramas. We then generate multiple perspective views from each panorama, following the systematic sampling strategy described in [97]. These views, along with the panoramas’ accurate GPS and pose data from large-scale SfM pipelines [47] that Google Street View is based on, serve as the high-quality georeferencing anchor for our 3D reconstructions.
@@
Feature Matching and SfM. To reconstruct from this heterogeneous set of smartphone and Street View images, we must establish robust feature matches. As a brute-force all-pairs matching approach is computationally infeasible, we adopt a retrieval-based strategy. We first compute a global descriptor for every image using EigenPlaces [2]. We then use these features to find the top 20 nearest neighbors for each image using the Faiss library [21], efficiently identifying pairs with likely visual overlap. For these candidate pairs, we perform local feature matching by extracting keypoints and descriptors with SuperPoint [19] and matching them with LightGlue [58]. With this graph of matched images, we perform Structure-from-Motion (SfM) using the global solver GLOMAP [75] to recover camera poses and a sparse 3D point cloud. COLMAP [85] is also applicable but GLOMAP [75] is an order-of-magnitude faster.
@@
Georeferencing and Trajectory Generation. A key step is georeferencing the resulting 3D reconstruction such that we align the reconstruction to a real-world coordinate system via a 7-DoF similarity transformation. Following the procedure described in [97, 96], this is computed by minimizing the discrepancy between the recovered poses of the Street View images and their ground-truth GPS data, which we project into an Earth-Centered, Earth-Fixed (ECEF) global coordinate frame. We explicitly discard the noisy GPS from the smartphone videos during this alignment, relying solely on the high-quality Street View data for metric accuracy [47]. The result is a single, georeferenced sparse reconstruction where the poses for all smartphone images are accurately localized, forming precise 3D trajectories. We then fit a ground plane to the reconstruction by using a Mask2Former [12] semantic segmentation model to identify 3D points corresponding to the road surface. By projecting the 3D camera poses onto this fitted plane, we define the vehicle’s 3D path, which is then projected back into the source images to create 2D drivable trajectories. Visualization of our trajectories can be viewed in Figure 14.
@@
Figure 14: Metric Geo-referenced Trajectories from our 3D Reconstruction Pipeline. We show examples of trajectories obtained from our reconstruction pipeline overlayed on birds-eye-view maps retrieved from OpenStreetMaps. We show some interesting situations for planning which require all aspects of scene perception, the trajectories are shown for a 4 second future horizon. (a-c) Depict a construction zone with two lane changes, first lane change to the right is marked by a TTC sign, while the second lane change is marked with drums. (d-e) Depict a construction zone marked with TTC signs and Vertical Panels. While there exists “free” space to navigate to the right most lane (where the workers are), the objects helpfully mark the actual drivable regions. Identifying drivable regions is still challenging for self-driving cars [22, 48]. (f) Cones behind a work zone vehicle mark it as a static object blocking the lane. This cue guides the cars to change the lane towards oncoming traffic to pass this work zone vehicle.
@@ C.1 Recognizing Work Zones
Faster R-CNN [80] 25.0 42.4 25.8 12.7 30.6 36.0
DiffusionDet [9] 31.1 50.1 32.2 18.3 30.8 42.0
Grounding DINO [61] 37.9 54.2 39.8 21.7 39.0 51.9
DINO [110] 39.9 57.2 42.2 24.0 38.6 52.1
Table 10: Detecting Work Zone Objects. We train detection models [9, 110] on the ROADWork dataset using a coarse vocabulary. The open-vocabulary detector [61] struggle to recognize work zone objects, but incorporating our data significantly improves its performance. Overall, our supervised models achieve substantially better results (+33.3 $AP$).
Detecting Work Zone Objects. As mentioned in Section 3 of the main manuscript, open-vocabulary detectors such as Grounding DINO [61] follow similar trends to Detic [115] and OpenSeeD [105]. As shown in Table 10, supervised models like DiffusionDet [9] and DINO [110] significantly outperform Grounding DINO (+33.3 $AP$). Fortunately, fine-tuning Grounding DINO [61] on our data almost matches the performance to DINO [110]. However, DINO [110] is still better than Grounding DINO [61] by +2.0 $AP$, likely reflecting the trade-off between a specialized detector and an open-vocabulary detector that generalizes well across a larger number of categories.
Method $AP$ $AP50$ $AP75$ $AP_{s}$ $AP_{m}$ $AP_{l}$
@@
Grounding DINO [61] (pre-trained) 5.1 7.9 5.3 2.2 5.3 9.1
DiffusionDet [9] 13.1 24.3 12.5 5.2 12.6 23.4
DINO [110] 19.7 32.1 20.2 10.0 18.0 31.8
Grounding DINO [61] 22.8 35.2 23.2 7.9 19.6 37.6
@@
Grounding DINO [61] (pre-trained) 8.8 13.0 9.5 6.6 10.9 11.7
DiffusionDet [9] 18.5 33.2 17.9 12.1 20.7 27.9
DINO [110] 27.3 43.0 28.2 17.0 28.8 36.2
Grounding DINO [61] 28.5 43.2 29.4 20.1 31.8 38.6
Table 11: Zero-Shot Detection On Discovered Workzones From BDD100K And Mapillary. For discovered-in-the-wild work zone images, fine-tuning the open-vocabulary detector Grounding DINO on our ROADWork dataset improves performance by +17.7 $AP$ on Mapillary and +19.7 $AP$ on BDD100K. Additionally, the supervised detectors DiffusionDet [9] and DINO [110] achieve promising performance.
Zero-Shot Detection On Discovered Workzones. In Section 3 of the main manuscript, we discovered 969 images in BDD100K [107] and Mapillary [74] datasets. While detectors trained on the ROADWork dataset facilitated the discovery of work zones around the world, their performance on these in-the-wild images has not been evaluated. To assess generalization, we manually annotated the 969 in-the-wild work zone images discovered in BDD and Mapillary (See Table 4 of the main manuscript for workzone discovery experiments). As shown in Table 11, the open-vocabulary detector Grounding DINO achieves significantly better zero-shot performance after being fine-tuned on our dataset, while supervised detectors also delivers promising zero-shot performance. Interestingly, compared to in-distribution performance (Table 10) where DINO [110] is better than Grounding DINO [61] by +2.0 $AP$, in this case Grounding DINO [61] shows improved generalization (+3.1 $AP$) over DINO [110].
Method $AP$ $AP50$ $AP75$ $AP_{s}$ $AP_{m}$ $AP_{l}$
@@
Detic [115] (pre-trained) 2.9 4.5 2.9 0.6 3.3 5.4
Mask R-CNN [33] 14.4 25.4 14.2 2.8 14.1 26.9
Mask DINO [51] 21.6 35.5 22.5 6.9 19 37.2
@@
Detic [115] (pre-trained) 3.7 5.8 4 3 5.1 4
Mask R-CNN [33] 19.8 33.8 21 12.8 23.3 28.1
Mask DINO [51] 29.1 46.6 31.3 18.1 31.4 45.5
Table 12: Zero-Shot Instance Segmentation on Discovered Workzones from BDD100K and Mapillary. As we noted in Section3, Detic [115] performed miserably for discovering work zones. We also observe that Detic’s zero-shot performance on work zone images from Mapillary [74] and BDD100K [107] follows the trends from the main manuscript. Similarly, Mask DINO [51] performs significantly better on both out-of-distribution datasets.
Zero-Shot Segmentation on Discovered Workzones. We evaluate open-vocabulary detectors and ROADWork supervised models on discovered images (See Section3) – which are out-of-distribution for all the models. Pre-trained Detic performs poorly on both Mapillary (2.9 AP) and BDD100K (3.7 AP), reinforcing our observation that work zone objects are severely underrepresented in foundation model training data. In contrast, models trained on the ROADWork dataset show substantial improvements. Mask DINO [51] achieves 21.6 AP on Mapillary and 29.1 AP on BDD100K, representing gains of +18.7 AP and +25.4 AP respectively over pre-trained Detic. Even the simpler Mask R-CNN architecture demonstrates significant improvements when trained on our dataset.
Method $AP$ $AP75$ $AP$ $AP75$ $AP$ $AP75$ $AP$ $AP75$
@@
Zero-Shot Instance Segmentation Results On Other Discovered Work Zone Images. We further evaluate our models on additional discovered work zone subsets (Vehicle-Pittsburgh, Vehicle-Rural, and Bus-Pittsburgh) to assess generalization under varying conditions. As shown in Table 13 pre-trained Detic [115] performs miserably across all subsets, with AP values ranging from 2.4 to 5.1. This performance is particularly poor in the Vehicle-Rural subset, where the AP is merely 2.4, highlighting the difficulty of segmenting work zones in rural environments. In contrast, models trained on ROADWork show substantially better performance, with Mask DINO achieving the best results across all subsets (+32.9 $AP$ on Vehicle-Pittsburgh, +27.7 $AP$ on Vehicle-Rural, +25.8 $AP$ on Bus-Pittsburgh Front Camera, and +28.9 $AP$ on Bus-Pittsburgh Side Camera compared to pre-trained Detic). These results further validate our observations from the main manuscript, confirming that foundation models struggle with work zone recognition in diverse conditions, while our ROADWork -trained models generalize effectively across various scenarios and viewpoints.
Training Data Test Data $AP$ $AP50$ $AP75$ $AP_{s}$ $AP_{m}$ $AP_{l}$
@@
Table 14: Does fine-tuning a open-vocabulary foundation model on ROADWork cause overfitting? Large foundation Models are prone to overfitting when trained on small datasets. To assess whether our dataset is large enough to mitigate overfitting, we finetune Grounding DINO [61] and evaluate it on common driving datasets including Cityscapes [17] and BDD100K [107], using their respective categories. We observe that fine-tuning significantly improves performance on ROADWork dataset (+31.3 $AP$), while performance on Cityscapes (+0.2 $AP$) and BDD100K (+0.1 $AP$) does not degrade.
Does fine-tuning a open-vocabulary foundation model on ROADWork cause overfitting? Large-scale foundation open-vocabulary model trained on millions of images may forget previously learned distributions when fine-tuned on additional data [40]. This effect is more pronounced when the dataset used for fine-tuning is small, leading to overfitting on the target data. To assess whether if ROADWork dataset is large enough to mitigate overfitting, we evaluate Grounding DINO [61] on common driving datasets Cityscapes [17] and BDD100K [107] with their label set as the vocabulary. We then finetune the model on ROADWork , and re-evaluate the detector on the same datasets with their vocabulary. As shown in Table 14, fine-tuning significantly improves performance on ROADWork (+31.3 $AP$), while performance marginally improves on Cityscapes [17] (+0.2 $AP$) or BDD100K [107] ($AP$). We hypothesize that this is due to the small domain gap between ROADWork dataset and common driving datasets. Hence, image features might remain consistent even when fine-tuned on our data. We leave further exploration of this phenomenon for future work.
Supervision $AP$ $AP50$ $AP75$ $AP_{s}$ $AP_{m}$ $AP_{l}$
Psuedo-Segmentations from SAM [46]
@@
Table 15: Are Manual Segmentations Still Needed? Results with Boundary IOU. We train instance segmentation models [33] with varying levels of supervision from the ROADWork dataset using SAM [46]. Unlike the results in the main paper, we evaluate performance using Boundary IOU [11], a metric more sensitive to boundary errors than standard IOU. We observe a larger improvement using Boundary IOU at higher thresholds (+6.7 $AP75_{IOU(B)}$), which indicates boundary quality improvements with manual annotations compared to psuedo ground truth annotations.
Are Manual Segmentations Still Needed? We posited in Section 3 of the main manuscript that some of the object categories in ROADWork exhibit irregular shapes. We present additional results in Table 15, computing AP using the Boundary IOU [11] metric ($AP_{IOU(B)}$). This metric penalizes boundary errors more strictly, making it more suitable for evaluating segmentation quality, particularly for irregularly shaped work zone objects such as arrow boards and work vehicles (e.g., “cranes”). Compared to the results in Section 3 of the main manuscript, we find that the performance gap at tighter thresholds is even more pronounced (+6.7 $AP75_{IOU(B)}$) compared to +5.3 $AP75$ from Table 3 of the main manuscript. This further underscores that manual ground-truth masks yield higher quality boundaries than those predicted by SAM [46], reinforcing the need for manual segmentations of rare work zone objects.

### v3

@@ B.1 Image Acquisition
Visual data were acquired from cameras mounted inside a vehicle while driving through 18 US cities, resulting in 9650 images from three sources: (a) images that we captured in Pittsburgh (b) images that were semi-automatically extracted from the Michelin Mobility Intelligence (MMI) Open Dataset (formerly RoadBotics) [roadbotics-open] (c) Images that were discovered in Mapillary [neuhold2017mapillary], BDD100K [yu2020bdd100k] and other other data sources by our models trained on data from the first two sources.
Main Data Sources. To collect the first data subset of the main dataset, we drove on urban, suburban, and rural roads in Pittsburgh and captured 2,338 ($32\%$) images with an iPhone 14 Pro Max paired with a Bluetooth remote trigger. Next, images from other U.S. cities were sourced from videos in the MMI Open Dataset. A combination of Detic [zhou2022detecting] and a cone detector trained on NuScenes [nuscenes] were used to mine frames presumed to contain roadwork zones with detector confidence at $25\%$ – ensuring high recall with the expense of low precision. This process yielded approximately 100000 candidate images. We then manually selected 5078 (68%) images containing unique road objects or roadwork zones, prioritizing individual scene diversity. The distribution of images across U.S. cities is shown in Figure 12.
Discovered Data Sources. In Section 3 of the main manuscript, we described our model and the work zone classification rule that we use to discover images. We discovered work zone images from common driving datasets (a) 558 images from Mappilary [neuhold2017mapillary] and (b) 411 images from BDD100K [yu2020bdd100k]. Additionally, we exploit other data sources to curate 1265 images into various subsets containing work zones (See Figure 13 for examples). These subsets were further manually filtered to remove redundant images. We describe the subsets below,
@@
Figure 13: ROADWork: Additional In-The-Wild Discovered and Annotated Work Zone Images. Following the discovery process described in Section 3 of the main manuscript, we discovered and annotated an additional 1265 work zones images from a variety of sources apart from 969 images discovered in BDD100K [yu2020bdd100k] and Mapillary [neuhold2017mapillary]. (a) The top row depicts Vehicle-Pittsburgh subset images we discovered from driving in Pittsburgh (around 157 images). The subset consists of work zone images taken in bad weather and night. (b) The middle row depicts Vehicle-Rural subset images we driving on rural areas and highways in the US (308 images). The subset consists of work zone images taken in both day and night. (c) The bottom row depicts images discovered from a Bus that was driven on a fixed route in Pittsburgh. We discovered 272 images captured from the front camera, while 528 images were captured from other cameras mounted on the bus. Images were captured in all conditions, including bad weather and night.
@@ B.2 Annotations
The object annotation workflow combined automatic and manual labeling, followed by manual verification. To reduce annotation effort, we used Detic [zhou2022detecting] with a custom vocabulary of “cone, drum, vehicle, traffic sign” to bootstrap annotations on our captured images. However, category predictions from Detic [zhou2022detecting] were discarded as due to frequent classification errors. Polygons were simplified using the Vishwalingam-Wyatt algorithm [visvalingam2017line] to facilitate editing. All object categories were manually assigned, and any additional objects in these images, as well as objects in all other images, were manually segmented and categorized. Finally, all annotations were manually verified by one person.
@@ B.3 Metric 3D Reconstruction and Pathway Generation from Smartphone Videos
Leveraging Smartphone-As-Dashcam Videos. Our work utilizes the MMI Open Data Set [roadbotics-open], which contains extensive video footage captured from a Samsung Galaxy S9 smartphone, for which camera intrinsics are known. From this dataset, we extract 30-second video snippets corresponding to our annotated work zones. These snippets are then downsampled to 5 FPS to yield the final set of smartphone images for our 3D reconstruction pipeline.
@@
Leverging 3D Reconstruction As Anchor. Our primary goal is to produce an accurate, metric-scale, and spatially-aligned 3D reconstruction from the collection of smartphone videos, which have weakly-aligned GPS metadata. With recent advances in Visual Place Recognition [berton2025megaloc], it’s likely that the weakly-aligned GPS metadata might also be superfluous in the future.
@@
The core of our approach is to anchor our reconstruction to a set of images with high-quality pose information [vuong2024toward, vuong2025aerialmegadepth]. To achieve this, we use the initial, coarse GPS from each video to query and retrieve nearby Google Street View panoramas. We then generate multiple perspective views from each panorama, following the systematic sampling strategy described in [vuong2024toward]. These views, along with the panoramas’ accurate GPS and pose data from large-scale SfM pipelines [klingner2013street] that Google Street View is based on, serve as the high-quality georeferencing anchor for our 3D reconstructions.
@@
Feature Matching and SfM. To reconstruct from this heterogeneous set of smartphone and Street View images, we must establish robust feature matches. As a brute-force all-pairs matching approach is computationally infeasible, we adopt a retrieval-based strategy. We first compute a global descriptor for every image using EigenPlaces [berton2023eigenplaces]. We then use these features to find the top 20 nearest neighbors for each image using the Faiss library [douze2024faiss], efficiently identifying pairs with likely visual overlap. For these candidate pairs, we perform local feature matching by extracting keypoints and descriptors with SuperPoint [detone2018superpoint] and matching them with LightGlue [lindenberger2023lightglue]. With this graph of matched images, we perform Structure-from-Motion (SfM) using the global solver GLOMAP [pan2024global] to recover camera poses and a sparse 3D point cloud. COLMAP [schoenberger2016sfm] is also applicable but GLOMAP [pan2024global] is an order-of-magnitude faster.
@@
Georeferencing and Trajectory Generation. A key step is georeferencing the resulting 3D reconstruction such that we align the reconstruction to a real-world coordinate system via a 7-DoF similarity transformation. Following the procedure described in [vuong2024toward, vuong2024walt3d], this is computed by minimizing the discrepancy between the recovered poses of the Street View images and their ground-truth GPS data, which we project into an Earth-Centered, Earth-Fixed (ECEF) global coordinate frame. We explicitly discard the noisy GPS from the smartphone videos during this alignment, relying solely on the high-quality Street View data for metric accuracy [klingner2013street]. The result is a single, georeferenced sparse reconstruction where the poses for all smartphone images are accurately localized, forming precise 3D trajectories. We then fit a ground plane to the reconstruction by using a Mask2Former [cheng2022masked] semantic segmentation model to identify 3D points corresponding to the road surface. By projecting the 3D camera poses onto this fitted plane, we define the vehicle’s 3D path, which is then projected back into the source images to create 2D drivable trajectories. Visualization of our trajectories can be viewed in Figure 14.
@@
Figure 14: Metric Geo-referenced Trajectories from our 3D Reconstruction Pipeline. We show examples of trajectories obtained from our reconstruction pipeline overlayed on birds-eye-view maps retrieved from OpenStreetMaps. We show some interesting situations for planning which require all aspects of scene perception, the trajectories are shown for a 4 second future horizon. (a-c) Depict a construction zone with two lane changes, first lane change to the right is marked by a TTC sign, while the second lane change is marked with drums. (d-e) Depict a construction zone marked with TTC signs and Vertical Panels. While there exists “free” space to navigate to the right most lane (where the workers are), the objects helpfully mark the actual drivable regions. Identifying drivable regions is still challenging for self-driving cars [news-2024missionlocal, news-nyt2023concrete]. (f) Cones behind a work zone vehicle mark it as a static object blocking the lane. This cue guides the cars to change the lane towards oncoming traffic to pass this work zone vehicle.
@@ C.1 Recognizing Work Zones
Faster R-CNN [ren2016faster] 25.0 42.4 25.8 12.7 30.6 36.0
DiffusionDet [chen2023diffusiondet] 31.1 50.1 32.2 18.3 30.8 42.0
Grounding DINO [liu2024grounding] 37.9 54.2 39.8 21.7 39.0 51.9
DINO [zhangdino] 39.9 57.2 42.2 24.0 38.6 52.1
Table 10: Detecting Work Zone Objects. We train detection models [chen2023diffusiondet, zhangdino] on the ROADWork dataset using a coarse vocabulary. The open-vocabulary detector [liu2024grounding] struggle to recognize work zone objects, but incorporating our data significantly improves its performance. Overall, our supervised models achieve substantially better results (+33.3 $AP$).
Detecting Work Zone Objects. As mentioned in Section 3 of the main manuscript, open-vocabulary detectors such as Grounding DINO [liu2024grounding] follow similar trends to Detic [zhou2022detecting] and OpenSeeD [xu2023open]. As shown in Table 10, supervised models like DiffusionDet [chen2023diffusiondet] and DINO [zhangdino] significantly outperform Grounding DINO (+33.3 $AP$). Fortunately, fine-tuning Grounding DINO [liu2024grounding] on our data almost matches the performance to DINO [zhangdino]. However, DINO [zhangdino] is still better than Grounding DINO [liu2024grounding] by +2.0 $AP$, likely reflecting the trade-off between a specialized detector and an open-vocabulary detector that generalizes well across a larger number of categories.
Method APAP $AP50$ $AP75$ $AP_{s}$ $AP_{m}$ $AP_{l}$
@@
Grounding DINO [liu2024grounding] (pre-trained) 5.1 7.9 5.3 2.2 5.3 9.1
DiffusionDet [chen2023diffusiondet] 13.1 24.3 12.5 5.2 12.6 23.4
DINO [zhangdino] 19.7 32.1 20.2 10.0 18.0 31.8
Grounding DINO [liu2024grounding] 22.8 35.2 23.2 7.9 19.6 37.6
@@
Grounding DINO [liu2024grounding] (pre-trained) 8.8 13.0 9.5 6.6 10.9 11.7
DiffusionDet [chen2023diffusiondet] 18.5 33.2 17.9 12.1 20.7 27.9
DINO [zhangdino] 27.3 43.0 28.2 17.0 28.8 36.2
Grounding DINO [liu2024grounding] 28.5 43.2 29.4 20.1 31.8 38.6
Table 11: Zero-Shot Detection On Discovered Workzones From BDD100K And Mapillary. For discovered-in-the-wild work zone images, fine-tuning the open-vocabulary detector Grounding DINO on our ROADWork dataset improves performance by +17.7 APAP on Mapillary and +19.7 APAP on BDD100K. Additionally, the supervised detectors DiffusionDet [chen2023diffusiondet] and DINO [zhangdino] achieve promising performance.
Zero-Shot Detection On Discovered Workzones. In Section 3 of the main manuscript, we discovered 969 images in BDD100K [yu2020bdd100k] and Mapillary [neuhold2017mapillary] datasets. While detectors trained on the ROADWork dataset facilitated the discovery of work zones around the world, their performance on these in-the-wild images has not been evaluated. To assess generalization, we manually annotated the 969 in-the-wild work zone images discovered in BDD and Mapillary (See Table 4 of the main manuscript for workzone discovery experiments). As shown in Table 11, the open-vocabulary detector Grounding DINO achieves significantly better zero-shot performance after being fine-tuned on our dataset, while supervised detectors also delivers promising zero-shot performance. Interestingly, compared to in-distribution performance (Table 10) where DINO [zhangdino] is better than Grounding DINO [liu2024grounding] by +2.0 $AP$, in this case Grounding DINO [liu2024grounding] shows improved generalization (+3.1 $AP$) over DINO [zhangdino].
Method APAP $AP50$ $AP75$ $AP_{s}$ $AP_{m}$ $AP_{l}$
@@
Detic [zhou2022detecting] (pre-trained) 2.9 4.5 2.9 0.6 3.3 5.4
Mask R-CNN [he2017mask] 14.4 25.4 14.2 2.8 14.1 26.9
Mask DINO [li2023maskdino] 21.6 35.5 22.5 6.9 19 37.2
@@
Detic [zhou2022detecting] (pre-trained) 3.7 5.8 4 3 5.1 4
Mask R-CNN [he2017mask] 19.8 33.8 21 12.8 23.3 28.1
Mask DINO [li2023maskdino] 29.1 46.6 31.3 18.1 31.4 45.5
Table 12: Zero-Shot Instance Segmentation on Discovered Workzones from BDD100K and Mapillary. As we noted in Section3, Detic [zhou2022detecting] performed miserably for discovering work zones. We also observe that Detic’s zero-shot performance on work zone images from Mapillary [neuhold2017mapillary] and BDD100K [yu2020bdd100k] follows the trends from the main manuscript. Similarly, Mask DINO [li2023maskdino] performs significantly better on both out-of-distribution datasets.
Zero-Shot Segmentation on Discovered Workzones. We evaluate open-vocabulary detectors and ROADWork supervised models on discovered images (See Section3) – which are out-of-distribution for all the models. Pre-trained Detic performs poorly on both Mapillary (2.9 AP) and BDD100K (3.7 AP), reinforcing our observation that work zone objects are severely underrepresented in foundation model training data. In contrast, models trained on the ROADWork dataset show substantial improvements. Mask DINO [li2023maskdino] achieves 21.6 AP on Mapillary and 29.1 AP on BDD100K, representing gains of +18.7 AP and +25.4 AP respectively over pre-trained Detic. Even the simpler Mask R-CNN architecture demonstrates significant improvements when trained on our dataset.
Method APAP $AP75$ APAP $AP75$ APAP $AP75$ APAP $AP75$
@@
Zero-Shot Instance Segmentation Results On Other Discovered Work Zone Images. We further evaluate our models on additional discovered work zone subsets (Vehicle-Pittsburgh, Vehicle-Rural, and Bus-Pittsburgh) to assess generalization under varying conditions. As shown in Table 13 pre-trained Detic [zhou2022detecting] performs miserably across all subsets, with AP values ranging from 2.4 to 5.1. This performance is particularly poor in the Vehicle-Rural subset, where the AP is merely 2.4, highlighting the difficulty of segmenting work zones in rural environments. In contrast, models trained on ROADWork show substantially better performance, with Mask DINO achieving the best results across all subsets (+32.9 $AP$ on Vehicle-Pittsburgh, +27.7 $AP$ on Vehicle-Rural, +25.8 $AP$ on Bus-Pittsburgh Front Camera, and +28.9 $AP$ on Bus-Pittsburgh Side Camera compared to pre-trained Detic). These results further validate our observations from the main manuscript, confirming that foundation models struggle with work zone recognition in diverse conditions, while our ROADWork -trained models generalize effectively across various scenarios and viewpoints.
Training Data Test Data APAP $AP50$ $AP75$ $AP_{s}$ $AP_{m}$ $AP_{l}$
@@
Table 14: Does fine-tuning a open-vocabulary foundation model on ROADWork cause overfitting? Large foundation Models are prone to overfitting when trained on small datasets. To assess whether our dataset is large enough to mitigate overfitting, we finetune Grounding DINO [liu2024grounding] and evaluate it on common driving datasets including Cityscapes [cordts2016cityscapes] and BDD100K [yu2020bdd100k], using their respective categories. We observe that fine-tuning significantly improves performance on ROADWork dataset (+31.3 $AP$), while performance on Cityscapes (+0.2 $AP$) and BDD100K (+0.1 APAP) does not degrade.
Does fine-tuning a open-vocabulary foundation model on ROADWork cause overfitting? Large-scale foundation open-vocabulary model trained on millions of images may forget previously learned distributions when fine-tuned on additional data [ilharco2022patching]. This effect is more pronounced when the dataset used for fine-tuning is small, leading to overfitting on the target data. To assess whether if ROADWork dataset is large enough to mitigate overfitting, we evaluate Grounding DINO [liu2024grounding] on common driving datasets Cityscapes [cordts2016cityscapes] and BDD100K [yu2020bdd100k] with their label set as the vocabulary. We then finetune the model on ROADWork , and re-evaluate the detector on the same datasets with their vocabulary. As shown in Table 14, fine-tuning significantly improves performance on ROADWork (+31.3 $AP$), while performance marginally improves on Cityscapes [cordts2016cityscapes] (+0.2 $AP$) or BDD100K [yu2020bdd100k] ($AP$). We hypothesize that this is due to the small domain gap between ROADWork dataset and common driving datasets. Hence, image features might remain consistent even when fine-tuned on our data. We leave further exploration of this phenomenon for future work.
Supervision APAP $AP50$ $AP75$ $AP_{s}$ $AP_{m}$ $AP_{l}$
Psuedo-Segmentations from SAM [kirillov2023segany]
@@
Table 15: Are Manual Segmentations Still Needed? Results with Boundary IOU. We train instance segmentation models [he2017mask] with varying levels of supervision from the ROADWork dataset using SAM [kirillov2023segany]. Unlike the results in the main paper, we evaluate performance using Boundary IOU [cheng2021boundary], a metric more sensitive to boundary errors than standard IOU. We observe a larger improvement using Boundary IOU at higher thresholds (+6.7 $AP75_{IOU(B)}$), which indicates boundary quality improvements with manual annotations compared to psuedo ground truth annotations.
Are Manual Segmentations Still Needed? We posited in Section 3 of the main manuscript that some of the object categories in ROADWork exhibit irregular shapes. We present additional results in Table 15, computing AP using the Boundary IOU [cheng2021boundary] metric ($AP_{IOU(B)}$). This metric penalizes boundary errors more strictly, making it more suitable for evaluating segmentation quality, particularly for irregularly shaped work zone objects such as arrow boards and work vehicles (e.g., “cranes”). Compared to the results in Section 3 of the main manuscript, we find that the performance gap at tighter thresholds is even more pronounced (+6.7 $AP75_{IOU(B)}$) compared to +5.3 $AP75$ from Table 3 of the main manuscript. This further underscores that manual ground-truth masks yield higher quality boundaries than those predicted by SAM [kirillov2023segany], reinforcing the need for manual segmentations of rare work zone objects.